Demand curve: P = 30 – Q Supply curve: P = 2Q

  1. a) Calculate the equilibrium quantity and price.
  2. b) Draw the curves.
  3. c) Suppose that the government set the price at 25 dollars. Calculate the shortage or

surplus that is created on the market.

W4 exercise

Use the demand and supply functions from ‘W3 exercise’ and calculate the consumer surplus and the producer surplus.

W5 exercise

Suppose that the demand schedule for electric bicycles is as follows:

  1. a) Use the arc price elasticity calculation method to calculate the price elasticity of demand as the price increases from CHF1,500 to 2,000 if
    1. the income level is CHF70,000
    2. the income level is CHF100,000
  2. b) Calculate the income elasticity of demand as the income increases from CHF70,000 to CHF100,000 if the price is set at CHF2,000.
Answer #1

Answers to Question W2 and W3

- 404 Price Supply Price Supply With no price regulation 300 (0,30) Consumer Surplus P=$25 price set by Govt (10,20) (10, 20)

Demand curve: P = 30 – Q or Qd = 30-P and the Supply curve: P = 2Q or Qs = P/2

Equilibrium at the intersection of demand and supply curves, so we equate the two equations to get equilibrium quantity and substitute the value in any of the above given function to arrive at the equilibrium price.

30 – Q = 2Q

3Q = 30 or Q = 10

at Q =10 P = 2*10 = 20

The price set by the government is $25.

At P = $25, Quantity demanded or Qd = 30-25 = 5

At P = $25, Quantity supplied or Qs = 25/2 = 12.5

Thus there is excess supply in the market of 12.5-5 = 7.5 units

Consumer Surplus if there is NO price regulation = 0.5*(30-20)*10 = 50

Producer Surplus if there is NO price regulation = 0.5*(20-0)*10 = 100

Consumer Surplus if there is price regulation = 0.5*(30-25)*5 = 12.5

Producer Surplus if there is no price regulation = 0.5*(10-0)*5+(25-10)*5 = 100